Calculate the pH of a buffer that is 0.012 M C6H5COOH ( Ka= 6.3x10-5 ) and 0.033 M NaC6H5COO.


1
Expert's answer
2021-06-21T06:47:14-0400


Solution


pH=pKa+log"\frac{[A^-]}{[HA]}"


pKa=-logKa=-log("6.3\times10^{-5})=4.20"

[A-]=0.033M

[HA]=0.012M


pH=4.20+log"(\frac{0.033}{0.012})"


pH= 4.639

pH=4.64